Princeton University: Plasma Physics Laboratory

To:

Dennis O’Brien Lang Lao Distribution

From:

Dick Wieland Doug McCune Date: February 3, 1997

Subject: EFIT and TRANSP Collaboration

Here is a long promised attempt at summarizing our recent discussions on a proposed collaboration to merge EFIT into TRANSP. Please consider this a draft statement, and feel free to modify it as you see fit. If we can produce a mutually satisfactory statement, then probably the next step would be to pass this on to our respective management for further action.

In discussions here at PPPL we’ve agreed that a free-boundary equilibrium should be the next major upgrade to TRANSP. It will provide us with the tools we need to better analyze TFTR and PBX-M plasmas, as well as making TRANSP better able to analyze your JET and DIII-D divertor plasmas.

We’ve been considering which free boundary equilibrium code(s) to use in TRANSP, and conclude that EFIT is the best suited of the “direct” Cartesian codes for this purpose. We will probably also try to use an “inverse” moments code as well, but haven’t yet determined which one.

We’ve talked amongst ourselves informally about establishing some sort of collaboration to incorporate the EFIT code into TRANSP. We at PPPL think that this is a great idea, and recommend that we try to start something up at the beginning of the new year.

The proposed upgrade to TRANSP will be extensive and require work in several areas:

TRANSP

• Expand the input data structure to accomodate data needed to run EFIT

• Expand the output data structure and RPLOT to provide additional display information related to EFIT.

• Expand the TRANSP architecture to accomodate external measurements of I(R), either directly or through the equilibrium solver.

• Construct an interface between the “inverse” TRANSP flux surface geometry {R, Z}(ρ, θ) and the “direct” EFIT geometry Ψ(R, Z).

• Construct a TRANSP-based time sequence equilibrium driver for studying the timedependent behavior of free boundary equilibrium codes with TRDAT filtered input data.

EFIT

• As a first step, we would like to familiarize ourselves here at PPPL with EFIT, by setting up a local version and running it stand alone on TFTR and PBX-M plasmas.

February 3, 1997 Page 2

• Construct call interfaces for EFIT so that it can be called from within TRANSP, and reorganize code to facilitate this new mode of operation.

• Construct error handling and self-correction algorithms to withstand the inhospitable profile environment that TRANSP can present

TRANSP ⇔ EFIT interplay

• Time will be required to fully exercise the TRANSP-EFIT interface and work out the wrinkles.

We could proceed with these topics in such a way that PPPL and JET collaborate on the TRANSP-EFIT interface, with GA providing consultative support on EFIT equilibrium and coding issues.

A proposed timetable for the project might look like this:

1st & 2nd Quarters 1995 PPPL work on expanding TRANSP data structures and logic to accomodate freeboundary mode of operation; gain local expertise with EFIT. JET work on subroutinization of EFIT and planning its interface to TRANSP.

3rd Quarter 1995 PPPL and JET collaborate on implementing the TRANSP-EFIT interface and debugging the code.

We should also propose to management that funds be made available for short term and long term travel, so that periodic exchange visits could be arranged to facilitate achieving the objectives of the program.

Distribution: Mike Zarnstorff